OpenStudy (anonymous):

consider points A(-3,5) and point B(x,2x). Find point B so that the distance between A and B is exactly √37.

OpenStudy (anonymous):

Do you know the distance formula between 2 points... or the Pythagorean theorem for right triangles?

OpenStudy (anonymous):

i tried the distance formula and set it equal to √37

OpenStudy (anonymous):

that's a good approach... how did you set up the distance formula? Maybe you just made a small error...

OpenStudy (anonymous):

i got down to 4x^2-13x-53=37

OpenStudy (anonymous):

i set it up like √37= √(-3x-x)^2-(5-2x)^2

OpenStudy (anonymous):

That first parenthesis should be (-3-x)^2 not (-3x-x)^2

OpenStudy (anonymous):

yea sorry that was just a typo

OpenStudy (anonymous):

stuff inside the square root: = (-3-x)^2 - (5-2x)^2 = 9-6x+x^2 - 25 -20x + 4x^2 = 5x^2 -26x -16

OpenStudy (anonymous):

shouldn't the -6x on the second line be +6x

OpenStudy (anonymous):

yes, you're right.. .good catch, sorry...

OpenStudy (anonymous):

so it should be 5x^2 -14x -16 ?

OpenStudy (anonymous):

yes. i found my mistake of not having the single "x" squared. i forgot to square it and left it only as x

OpenStudy (anonymous):

ok :) It's a pain when you make a small error like that... takes awhile to find it sometimes. Did you already get a solution for x?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

did you find an answer i am having trouble

OpenStudy (anonymous):

Sorry for the delay... let me work it through from the beginning to make sure I don't keep injecting dumb errors due to working to quickly

OpenStudy (anonymous):

no problem. Thanks

OpenStudy (anonymous):

i got down to 5x^2-14x-53. and i was stuck. Is this atleast right?

OpenStudy (anonymous):

Sqrt(37) = Distance between points = \[\sqrt{5x^2 - 14x + 34}\]

OpenStudy (anonymous):

how did you get +34?

OpenStudy (anonymous):

I didn't, yet... but that should be the expression for distance between the 2 points. I was having browser issues... didn't want to hold you up if you were ahead of me.

OpenStudy (anonymous):

oo okay

OpenStudy (anonymous):

(x- (-3))^2 = (x+3)^2 = x^2 + 6x +9 (2x-5)^2 = 4x^2 -20x + 25 inside sqrt: (x^2 + 6x +9) + (4x^2 -20x + 25) = 5x^2 -14x + 34 all that is = sqrt(37), so... sqrt(37) = sqrt( 5x^2 -14x +34) 0 = 5x^2 -14x +34 -37 = 5x^2 - 14x -3 0 = (5x -1)(x - 3) x = -1/5 OR x = 3

OpenStudy (anonymous):

Yea that is what I just got too Thanks

OpenStudy (anonymous):

thanks for your patience :) Looks straightforward to just use "distance formula", but it's tricky trying not to make algebra errors along the way.

OpenStudy (anonymous):

tell me about it

OpenStudy (anonymous):

one last thing

OpenStudy (anonymous):

to answer the question would you plug in both points to see which one works

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yeah, I was thinking about that too... My instinct tells me that since both are valid solutions for x, there are in fact 2 points (x,2x) in which the distance from the other point is sqrt(37).

OpenStudy (anonymous):

You could confirm by substituting both, but they should BOTH work.

OpenStudy (anonymous):

so my answer would be both... x=3 AND -1/5?

OpenStudy (anonymous):



OpenStudy (anonymous):

What exactly does this mean?

OpenStudy (anonymous):

it means two different points satisfy these two constraints: Constraint 1: solution point has distance sqrt(37) from the point (-3,5) Constraint 2: y value of the point = twice x value of the point i.e. (x,2x)

OpenStudy (anonymous):

sorry that was a typo

OpenStudy (anonymous):

my answer for point be would b both (3,6) and (-1/5, -2/5)??

OpenStudy (anonymous):

yes, that's it. It seems weird, I guess, but there are really two valid solutions for the point described in the problem... and if my drawing had been more to scale, you would see it easier there.

OpenStudy (anonymous):

ok Thank you

OpenStudy (anonymous):

glad to help :)